# Task-free Lifelong Robot Learning with Retrieval-based Weighted Local Adaptation

### v3

@@ A.3.2 Description Paraphrase
Figure 8: Task Blurry Effect on libero_spatial benchmark. After paraphrasing the task descriptions, both BERT and SS models struggle to distinguish the tasks in libero_spatial.
@@
A.4 Details about Selective Weighting
In this section, we introduce our Selective Weighting mechanism in detail.
A.4.1 Detailed Heuristics and Implementations
To assign weights to retrieved demonstrations, we analyze the image embedding distance between demonstration and failed rollout trajectories. Typically, the embedding distance increases as the failed rollout diverges from the demonstration. We selectively add weights for the frames in the retrieved demonstration using the Embedding Distance Curve (EDC), derived from the Embedding Distance Matrix (EDM), as illustrated in Figure 10.
Due to the multi-modal nature of robotic actions and visual observation noise, raw embedding distances can be erratic. To mitigate this, we smooth the EDC using a moving average window. Despite smoothing, the curve may remain jittery, making it difficult to pinpoint a single divergence point. Therefore, we identify a range of frames, termed the Separation Segment, where the distances increase, indicating vulnerable steps that lead to task failure.
We apply two thresholds to identify the segment: a lower threshold at $\frac{1}{8}$ and an upper threshold at $\frac{1}{3}$ of the maximum observed distance in EDC. We locate frames where the smoothed embedding distance falls within this range, focusing on the last occurrence to account for initial divergences that may later converge. We then extend this segment by $15$ frames before and after to mitigate noise effects.
For each frame within the Separation Segment, we increment the corresponding weight in the initially uniform weight vector by $0.3$. This process is repeated for up to five failed rollouts per retrieved demonstration. After processing all demonstrations, we clip the weights to a maximum of $2$ and normalize the weight vector to maintain consistent loss scaling and ensure stable gradient updates. During adaptation, the resulting weights ($w_{t,n}$) are integrated into the loss function as described in Equation (2). This selective weighting emphasizes critical samples while reducing the influence of less relevant ones, thereby enhancing the model’s learning efficiency.
A.4.2 Hyperparameter Sensitivity Analysis on Selective Weighting.
@@ A.4.1 Detailed Heuristics and Implementations
A.5 Detailed Testing Results
We selected 20 typical scenarios among libero_90. The list of those scenarios can be found in Table 7. Additionally, the testing results of our method and baselines including ER-RWLA, ER, Packnet, are listed in Table 8
A.6 Discussion on Potential Forgetting during Local Adaptation

### v4

@@ A.3.2 Description Paraphrase
Figure 8: Task Blurry Effect on libero_spatial benchmark. After paraphrasing the task descriptions, both BERT and $S$ models struggle to distinguish the tasks in libero_spatial.
@@
A.4 Detailed Heuristics and Implementations for Selective Weighting
Selective weighting strives to concentrate the adaptation loss on the critical steps of a demonstration—those at which the policy is most likely to diverge, forget, or fail. We detect these steps by measuring how a failed rollout departs from the demonstration in the image–embedding space.
For every retrieved demonstration we compute an Embedding Distance Matrix (EDM) between its frames and those of up to five failed rollouts. A row-wise minimum over the EDM yields the Embedding Distance Curve (EDC), which tracks, step-by-step, the smallest visual discrepancy seen so far. Because raw embedding distances are noisy (multi-modal actions, rendering noise), we smooth the EDC with a moving average. The model typically fails when the EDC rises and stays high. We mark as a Separation Segment the last contiguous block of frames whose smoothed distance is between $\frac{1}{8}$ and $\frac{1}{3}$ of the curve’s maximum. To further mitigate noise effects, we pad this block by $\pm 15$ frames.
Starting from a uniform weight vector, we add $0.3$ to every frame inside the Separation Segment (for each retrieved demonstration based on the failed rollouts). Then, we clip weights to 2 and $\ell_{1}$-normalise all weights in each demonstration, producing $w_{t,n}$ used in Eq. equation 3. This focuses gradient updates on the vulnerable steps while keeping the overall loss scale stable.
Specifically, the hyperparameters such as thresholds ($\frac{1}{8}$, $\frac{1}{3}$) and padding (15 frames) were tuned on libero_object to best capture true divergence points; performance gains generalize across libero_spatial and libero_goal (Table 2).
A.5 Additional Experiments
A.5.1 Comparison with Adaptation-based baseline
Our method can also be regarded as a form of test-time adaptation, akin to Schmied et al. (2023); Liu et al. (2023); Singh et al. (2024); Peng et al. (2020); de Masson D’Autume et al. (2019). To benchmark against an adaptation-oriented alternative, we chose LoRA-FT (Schmied et al., 2023). During the one-step adaptation phase, LoRA-FT (i) equips visual, language, and extra-modality encoders with LoRA adapters, (ii) inserts LoRA on the queries and values of every Transformer decoder layer, and (iii) fully fine-tunes the policy head.
A model with the identical architecture is first pre-trained for 50 epochs on the libero_90 suite (90 short-horizon tasks) in a multitask fashion. We then adapt LoRA-FT to unseen tasks from libero_spatial, libero_object, and libero_goal. Both pre-training and adaptation use random seeds $\{1,\,21,\,42\}$. Unlike our ER-RWLA method—which has already experienced a lifelong-learning phase and therefore reviews only a limited subset of demonstrations—LoRA-FT is granted access to the entire dataset of each target task during adaptation.
Figure 10: Average success rate (± std) over three seeds on each benchmark. Both methods adapt for 20 epochs; LoRA-FT is allowed to see all demonstrations of the target task, whereas our proposed ER-RWLA only uses a small subset from $\mathcal{M}$.
Figure 10 shows that our method consistently surpasses LoRA-FT across all three benchmarks. This aligns with the observation of Liu et al. (2024): pre-training on many short-horizon tasks can hurt downstream continual-learning or adaptation performance, and training from scratch may even outperform such pre-trained models. A likely cause is distributional mismatch—the 90 pre-training tasks differ substantially from the evaluation benchmarks—limiting positive transfer in LoRA-FT.
Task ER-RWLA ER Packnet
0 0.85 $\pm$ 0.08 0.50 $\pm$ 0.03 1.00 $\pm$ 0.00
1 0.13 $\pm$ 0.08 0.27 $\pm$ 0.06 0.83 $\pm$ 0.09
2 0.73 $\pm$ 0.09 0.72 $\pm$ 0.10 0.92 $\pm$ 0.02
3 0.40 $\pm$ 0.03 0.13 $\pm$ 0.02 0.17 $\pm$ 0.03
4 0.93 $\pm$ 0.04 0.72 $\pm$ 0.10 1.00 $\pm$ 0.00
5 1.00 $\pm$ 0.00 0.57 $\pm$ 0.16 1.00 $\pm$ 0.00
6 0.52 $\pm$ 0.04 0.52 $\pm$ 0.03 0.78 $\pm$ 0.04
7 0.82 $\pm$ 0.07 0.63 $\pm$ 0.09 0.88 $\pm$ 0.02
8 0.32 $\pm$ 0.07 0.23 $\pm$ 0.06 0.00 $\pm$ 0.00
9 0.48 $\pm$ 0.15 0.38 $\pm$ 0.12 0.00 $\pm$ 0.00
10 0.23 $\pm$ 0.06 0.03 $\pm$ 0.02 0.00 $\pm$ 0.00
11 0.20 $\pm$ 0.03 0.10 $\pm$ 0.06 0.00 $\pm$ 0.00
12 0.23 $\pm$ 0.09 0.13 $\pm$ 0.02 0.00 $\pm$ 0.00
13 0.67 $\pm$ 0.09 0.83 $\pm$ 0.04 0.00 $\pm$ 0.00
14 0.15 $\pm$ 0.03 0.13 $\pm$ 0.04 0.00 $\pm$ 0.00
15 0.68 $\pm$ 0.09 0.30 $\pm$ 0.08 0.00 $\pm$ 0.00
16 0.03 $\pm$ 0.03 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00
17 0.28 $\pm$ 0.08 0.02 $\pm$ 0.02 0.00 $\pm$ 0.00
18 0.10 $\pm$ 0.08 0.02 $\pm$ 0.02 0.00 $\pm$ 0.00
19 0.27 $\pm$ 0.16 0.58 $\pm$ 0.07 0.00 $\pm$ 0.00
Note that, though we include such comparison with continual inference-time adaptation methods such as LoRA-FT, to assess our approach’s performance relative to, our motivations and settings still differ greatly. Continual inference-time adaptation often aims to enable models to deal with continual domain transfer with streaming data, while our approach focuses on mitigating catastrophic forgetting through a brief, efficient "review" phase prior to policy deployment.
A.5.2 Hyperparameter Sensitivity Analysis on Selective Weighting.
@@ A.5.1 Comparison with Adaptation-based baseline
A.6 Detailed Testing Results
We selected 20 typical scenarios among libero_90. The list of those scenarios can be found in Table 8. Additionally, the testing results of our method and baselines including ER-RWLA, ER, Packnet, are listed in Table 7.
A.7 Further Discussions
A.7.1 Potential Forgetting during Local Adaptation
@@
A.7.2 Future Work beyond LIBERO
Current LIBERO benchmarks treat similar scenarios (e.g., variations in object categories, layouts, or goals) as distinct tasks. However, this does not reflect how humans learn continually — we tend to generalize across such variations. Thus, we plan to identify equivariant features among similar tasks to avoid redundant retraining or adaptation. This strategy will further mitigate catastrophic forgetting and improve the system’s plasticity in more complex lifelong robot manipulation tasks with the proposed RWLA algorithm.